# SpikeNAS: A Fast Memory-Aware Neural Architecture Search Framework for Spiking Neural Network-based Autonomous Agents

### v3

@@ VI Conclusion
We propose the SpikeNAS framework to quickly find an SNN architecture that provides high accuracy under a memory constraint from an autonomous mobile agent. SpikeNAS employs an analysis of the impact of operations on the accuracy, architecture enhancements, and a NAS that is guided by memory budget and consecutive searches across cells. The experimental results show that, SpikeNAS maintains the accuracy and improves the searching time as compared to the state-of-the-art with a smaller memory cost. For instance, our SpikeNAS improves the searching time by 4.4x and accuracy by 1.3% under 2M parameters for CIFAR100. In this manner, our SpikeNAS framework enables the efficient SNN deployments for autonomous mobile agents.
@@ Acknowledgment
This work has been supported in parts by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.

### v4

@@ 6 Conclusion
We propose the SpikeNAS framework to quickly find an SNN architecture that provides high accuracy under a memory constraint from the hardware platform of the embedded AI system. SpikeNAS employs an analysis of the impact of operations on the accuracy, architecture enhancements, and a fast search algorithm that is guided by memory budget and consecutive searches across cells. The experimental results show that, SpikeNAS maintains the accuracy and improves the searching time as compared to the state-of-the-art with a smaller memory cost. As a result, our SpikeNAS improves the searching time and maintains high accuracy compared to state-of-the-art while meeting the given memory budgets (e.g., 29x, 117x, and 3.7x faster search for CIFAR10, CIFAR100, and TinyImageNet200, respectively). In this manner, our SpikeNAS framework enables the efficient SNN deployments in embedded AI systems for diverse application use-cases.
@@ Acknowledgment
This work was partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.